OpenStudy (anonymous):

Use the distributive property to factor the expression below. 3x-9x2

OpenStudy (anonymous):

@skullpatrol can u help

OpenStudy (skullpatrol):

Factor out the common factor first.

OpenStudy (anonymous):

Let's simplify step-by-step. 3x−9x2 =3x+−9x2 Answer: =−9x2+3x

OpenStudy (anonymous):

hs offline

OpenStudy (skullpatrol):

$$\Huge ab-ac=a(b-c)$$ $$\Huge a=3x $$$$\Huge b=1 $$$$\Huge c=3x $$ $$\Huge 3x-9x^2=3x(1-3x)$$